# Self-Organizing Dual-Buffer Adaptive Clustering Experience Replay (SODASER) for Safe Reinforcement Learning in Optimal Control

### v1

@@ Introduction
16:  Compute the membership strength of $S_{old}$ in all existing clusters $\mathcal{C}_{j}$, $\mu_{\mathcal{C}_{j}}(S_{old})$, using (21).
17:  where $\mathbf{c}_{j}$ is the center and $\sigma_{j}^{2}$ is the standard deviation of cluster $\mathcal{C}_{j}$;
18:  if for all clusters $\max(\mu_{\mathcal{C}}(S_{old}))\leq\Gamma_{th}$ then
19:   Add a new cluster $\mathcal{C}_{\text{new}}$ in the Cluster Buffer $\mathcal{C}$ based on $S_{old}$; $\mathbf{c}_{new}=S_{old}$ and $\sigma_{new}=\sigma_{0}=0.02$;
20:  else
21:   Update the cluster $\mathcal{C}_{j}$ with the maximum membership $\mu_{\mathcal{C}}(S_{old})$ by adding $S_{old}$ to $\mathcal{C}_{j}$; $C_{j}=\frac{\mathcal{N}_{j}C_{j}+S_{old}}{\mathcal{N}_{j}+1}$, $\mathcal{N}_{j}=\mathcal{N}_{j}+1$;
22:  end if;
23:  Variances Amplification: For the cluster $\mathcal{C}_{j}$ that covers $S_{old}$, increase variance $\sigma_{j}\leftarrow\sigma_{j}\times(1+\beta)$;
24:  Variances Reduction (Forgetting Factor): For all clusters $\mathcal{C}_{k}$, reduce variance $\sigma_{k}\leftarrow\sigma_{k}\times\sigma_{0}(\frac{1}{\rho})(1-\frac{\mathcal{N}_{k}}{\sum(\mathcal{N}_{i})})$;
25:  Omitting Narrow Clusters:
26:  if for all clusters $\sigma_{\mathcal{C}_{k}}\leq\sigma_{th}$ then
27:   delete the cluster $\mathcal{C}_{k}$;
28:  end if
29:  Similar Cluster Merge:
30:  if for all clusters $\|C_{i}-C_{j}\|<\gamma\times\max(\sigma_{i},\sigma_{j})$ then
31:   merge the clusters $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$ into $\mathcal{C}_{merged}$; $C_{\text{merged}}=\frac{\mathcal{N}_{i}C_{i}+\mathcal{N}_{j}C_{i}}{\mathcal{N}_{i}+\mathcal{N}_{j}}$,  $\mathcal{N}_{merged}=(\mathcal{N}_{i}+\mathcal{N}_{j})$, and $\sigma_{merged}=\max(\sigma_{i},\sigma_{j})$;
32:  end if
33:  Generate mini-batches $\mathcal{B}$: by sampling from all clusters in $\mathcal{C}$, with priority on clusters containing more samples; $\mathcal{B}\leftarrow$ { Fast-Buffer, Center of Clusters in Slow-Buffer, Random Samples of Slow-Buffer Cluster according to the clusters size };
34:  while weight convergence is achieved: $\varphi(t)^{\top}|W_{t+1}-W_{t}|\leq\delta$ do
35:   Compute gradients $\nabla_{W}J(W)$ on the mini-batches:
$\nabla_{W}J(W)=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\nabla_{W}L(\mathbf{x}_{i},\mathbf{u}_{i})$
36:   Update first moment estimate $m_{t}$ on the mini-batches
@@
37:   Update second moment estimate $v_{t}$ on the mini-batches
@@
38:   Correct bias in moment estimates
@@
39:   Update weights $W$ using the Sophia optimizer:
@@
40:  end while
41:  Update $t\leftarrow t+1$;
42:end while;
43:Output and analyze results based on the updated state variables and control inputs over time;
@@
This section presents a thorough assessment of the proposed methodology. Initially, an analysis of the system is conducted in the absence of control mechanisms, utilizing parameter values as specified in Table 1. The resulting simulation outcomes for the system states are depicted in Figure 3 (upper panel). Subsequently, fixed control values are introduced, and the dynamic evolution of system state variables under the influence of our proposed approach starting from predefined initial conditions—is illustrated in Figure 3 (lower panel). To further validate the adaptability of the proposed method, we consider random initial conditions and evaluate its performance over 200 iterations. The results are depicted in Figure 4 and Figure 5, showing the spectrum of state variables and control signals, respectively. In these figures, the blue line represents the average values over the 200 iterations, providing insights into the consistency and effectiveness of the proposed approach.
@@
Figure 5: Spectral representation of control signals using the proposed approach, based on 200 simulation runs.
@@
In Scenario 1 ($f_{1}$), vaccination rates for pre-sexually active individuals ($w_{1},w_{2}$), sexually active individuals ($u_{1},u_{2}$), and the female screening rate ($\alpha$) were set to zero. Scenario 2 ($f_{2}$) focused on the vaccination rates of sexually active individuals ($u_{1},u_{2}$) alongside the screening rate ($\alpha$), emphasizing the impact of targeting actively transmitting populations in reducing HPV prevalence. Scenario 3 ($f_{3}$) isolated the vaccination rates for sexually active individuals ($u_{1},u_{2}$), assessing their individual contribution to transmission mitigation. Scenario 4 ($f_{4}$) concentrated solely on the screening rate ($\alpha$), highlighting the critical role of detection and early intervention strategies in controlling HPV spread. Lastly, Scenario 5 ($f_{5}$) utilized all control variables dynamically, demonstrating the effectiveness of a fully integrated control strategy in optimizing HPV intervention outcomes. Table 2 presents a summary of scenarios, including their corresponding objective functions and activated control variables.
@@
The performance results, summarized in Table 3, demonstrate that SODACER-Sophia consistently outperformed RER and CBER across all metrics. For instance, in Scenario 5, which leverages the full set of controls dynamically, SODACER-Sophia achieved the best convergence rate and highest sample efficiency, highlighting its ability to optimize control strategies effectively. Figure 6 further illustrates the average results across 200 runs, showing SODACER-Sophia’s superior convergence performance compared to RER and CBER. Additionally, Figure 7 provides a detailed comparison of the performance spectrum across 200 runs for the value of the cost function in Equation (28). The results reveal that SODACER-Sophia exhibited the narrowest gap between the worst and best cases, indicating robust and reliable performance. In contrast, RER and CBER showed greater variability, with mean convergence values of approximately 3.5 and 0.2, respectively.
@@
Additionally, to evaluate the effectiveness of control strategies for the HPV model, five distinct cost functions were formulated, each representing a specific control scenario. In “Scenario 1", the cost function incorporates the full set of controls in problem (28), including vaccination rates for pre-sexually active females and males ($w_{1}(t),w_{2}(t)$), vaccination rates for sexually active individuals ($u_{1}(t),u_{2}(t)$), and the screening rate for females ($\alpha(t)$). “Scenario 2" focuses on vaccination of pre-sexually active individuals, utilizing only $w_{1}(t)$ and $w_{2}(t)$ to assess the impact of early immunization on reducing HPV prevalence. “Scenario 3" targets vaccination of sexually active individuals by employing $u_{1}(t)$ and $u_{2}(t)$, highlighting the role of immunizing sexually active populations in controlling transmission. “Scenario 4" isolates the screening rate $\alpha(t)$ as the sole control, emphasizing the effectiveness of detection and early intervention strategies. Lastly, in “Scenario 5", all controls ($w_{1}(t),w_{2}(t),u_{1}(t),u_{2}(t),\alpha(t)$) are set to fixed values, providing a baseline for evaluating the dynamic optimization capabilities of the proposed methodology. These scenarios enable a systematic analysis of individual and combined control strategies to optimize HPV intervention.

### v2

@@ Introduction
16:  Compute the membership strength of $S_{old}$ in all existing clusters $C_{j}$, $\mu_{C_{j}}(S_{old})$, using (21).
17:  if for all clusters $\max(\mu_{C}(S_{old}))\leq\Gamma_{th}$ then
18:   Add a new cluster $C_{\text{new}}$ in the Cluster Buffer $\mathcal{C}$ based on $S_{old}$; $m_{new}=S_{old}$ and $\sigma_{new}=\sigma_{0}=0.02$;
19:  else
20:   Update the cluster $C_{j}$ with the maximum membership $\mu_{C}(S_{old})$ by adding $S_{old}$ to $C_{j}$; $m_{j}=\frac{\mathcal{N}_{j}m_{j}+S_{old}}{\mathcal{N}_{j}+1}$, $\mathcal{N}_{j}=\mathcal{N}_{j}+1$;
21:  end if;
22:  Variances Amplification: For the cluster $C_{j}$ that covers $S_{old}$, increase variance $\sigma_{j}\leftarrow\sigma_{j}\times(1+\beta)$;
23:  Variances Reduction (Forgetting Factor): For all clusters $C_{k}$, reduce variance $\sigma_{k}\leftarrow\sigma_{k}\times\sigma_{0}(\frac{1}{\rho})(1-\frac{\mathcal{N}_{k}}{\sum(\mathcal{N}_{i})})$;
24:  Omitting Narrow Clusters:
25:  if for all clusters $\sigma_{C_{k}}\leq\sigma_{th}$ then
26:   delete the cluster $C_{k}$;
27:  end if
28:  Similar Cluster Merge:
29:  if for all clusters $\|m_{i}-m_{j}\|<\gamma\times\max(\sigma_{i},\sigma_{j})$ then
30:   merge the clusters $C_{i}$ and $C_{j}$ into $C_{merged}$; $m_{\text{merged}}=\frac{\mathcal{N}_{i}m_{i}+\mathcal{N}_{j}m_{j}}{\mathcal{N}_{i}+\mathcal{N}_{j}}$,  $\mathcal{N}_{merged}=(\mathcal{N}_{i}+\mathcal{N}_{j})$, and $\sigma_{merged}=\max(\sigma_{i},\sigma_{j})$;
31:  end if
32:  Generate mini-batches $B$: by sampling from all clusters in $\mathcal{C}$, with priority on clusters containing more samples; $B\leftarrow$ { Fast-Buffer, Center of Clusters in Slow-Buffer, Random Samples of Slow-Buffer Cluster according to the clusters size, sample with importance weights $w_{i}=\frac{1}{P(i)\cdot|C|}$ };
33:  while weight convergence is achieved: $\varphi(t)^{\top}|W_{t+1}-W_{t}|\leq\delta$ do
34:   Compute gradients $\nabla_{W}J(W)$ on the mini-batches:
$\nabla_{W}J(W)=\frac{1}{|B|}\sum_{i\in B}w_{i}\nabla_{W}L(\mathbf{x}_{i},\mathbf{u}_{i})$
35:   Update first moment estimate $M_{t}$ on the mini-batches
@@
36:   Update second moment estimate $v_{t}$ on the mini-batches
@@
37:   Correct bias in moment estimates
@@
38:   Update weights $W$ using the Sophia optimizer:
@@
39:  end while
40:  Update $t\leftarrow t+1$;
41:end while;
42:Output and analyze results based on the updated state variables and control inputs over time;
Computational Complexity Analysis
To provide a transparent assessment of the practical efficiency of the proposed SODACER framework, we analyze the computational cost associated with its dual-buffer architecture.
Theoretical Complexity
The primary computational operations in SODACER and their associated complexities are as follows:
  • Fast-Buffer Operations: Insertion and sampling from the First-In-First-Out (FIFO) Fast-Buffer require $O(1)$ time per experience. The memory requirement scales as $O(N_{fast})$, where $N_{fast}$ is the buffer capacity.
  • Clustering in Slow-Buffer: Assigning a transferred sample to the nearest cluster centroid involves a distance calculation to all $K$ clusters in a $d$-dimensional space, resulting in $O(K\cdot d)$ time complexity. Memory usage for storing $K$ clusters is $O(K\cdot d)$.
  • Buffer Transfer and Maintenance: The periodic transfer of samples from the Fast-Buffer to the Slow-Buffer and the subsequent cluster update (mean recalculation) are $O(1)$ and $O(d)$ operations, respectively. The less frequent cluster merging operation has a worst-case complexity of $O(K^{2})$ but is executed only at scheduled intervals.
Overall, the per-training-step overhead introduced by SODACER, compared to a standard uniform replay buffer, is dominated by the $O(K\cdot d)$ clustering assignment. This is typically negligible relative to the cost of neural network forward/backward passes, which scale with the model size and mini-batch dimensions.
Empirical Performance
We measured the average training time per iteration for the HPV control problem. The results indicate that the dual-buffer management incurs a consistent and bounded overhead. The clustered representation of the Slow-Buffer reduces the effective memory footprint by an order of magnitude compared to storing an equivalent number of raw experiences, enabling longer history retention without a linear increase in memory cost.
Practical Trade-offs and Optimizations
The moderate increase in computational cost per iteration is strategically traded for significant improvements in learning metrics:
  • Sample Efficiency: Intelligent sampling from diverse clusters reduces the number of environment interactions needed for convergence.
  • Memory Efficiency: Experience compression via clustering allows maintaining a broader experience distribution.
  • Stability: The separation of recent and historical data helps mitigate the bias-variance trade-off.
To further minimize overhead, clustering updates are performed asynchronously, and the most expensive operations (e.g., full similarity matrix calculation for merging) are scheduled during idle periods in the training loop.
In summary, the SODACER framework is designed with computational efficiency in mind. The introduced overhead is linear, predictable, and justified by the substantial gains in learning performance and memory utilization, making it a practical solution for complex reinforcement learning tasks.
@@
The safe set for the states of the system (28) is defined as
$\bar{X}=\left\{(U_{f},I_{f},V_{f},I_{m},V_{m})\in L^{2}(0,T)\mid 0\leq U_{f}(t),I_{f}(t),V_{f}(t),I_{m}(t),V_{m}(t)\leq 1\right\}$ (29)
Additionally, the CBFs are formulated as follows:
$B(X)=\frac{1}{2}X^{2},\quad\forall X\in\{U_{f},I_{f},V_{f},I_{m},V_{m}\}$ (30)
$B(J_{cost})=\frac{1}{2}J_{cost}^{2}.$
To ensure safety, the objective function is augmented with CBF terms $B(X)$ and $B(J_{cost})$ as follows:
$\underset{U_{f},I_{f},V_{f},I_{m},V_{m}\in\bar{X}}{\text{Minimize}}\quad J_{\text{spread}}=\int_{0}^{T}A_{0}(I_{m}+I_{f}+U_{f})dt+B(X)+B(J_{cost}),$ (31)
where $B(X):\bar{X}\longrightarrow\mathbb{R}$ for $X\in\{U_{f},I_{f},V_{f},I_{m},V_{m}\}$, and $B(J_{cost}):\mathrm{F}\longrightarrow\mathbb{R}$ exhibit the properties of CBFs.
@@
This section presents a thorough assessment of the proposed methodology. Initially, an analysis of the system is conducted in the absence of control mechanisms, utilizing parameter values as specified in Table 2. The resulting simulation outcomes for the system states are depicted in Figure 3 (upper panel). Subsequently, fixed control values are introduced, and the dynamic evolution of system state variables under the influence of our proposed approach starting from predefined initial conditions–is illustrated in Figure 3 (lower panel). To further validate the adaptability of the proposed method, we consider random initial conditions and evaluate its performance over 200 iterations. The results are depicted in Figure 4 and Figure 5, showing the spectrum of state variables and control signals, respectively. In these figures, the blue line represents the average values over the 200 iterations, providing insights into the consistency and effectiveness of the proposed approach. The control signals in Fig. 5 do not converge to zero because the optimal policy for the HPV model involves sustained vaccination and screening efforts to balance incoming susceptible individuals and ongoing transmission. This reflects a practical, cost-effective strategy where maintaining a certain level of intervention is optimal over the planning horizon, rather than driving controls to zero.
To assess the individual contribution of each functional module in the SODACER–Sophia framework, we conducted a series of ablation experiments. Three variants were compared against the full framework: (i) a single‑buffer configuration (using only the Fast‑Buffer), (ii) a Slow‑Buffer without adaptive clustering (replaced by a simple reservoir buffer), and (iii) replacement of the Sophia optimizer with the standard Adam optimizer.
The quantitative results are summarized in Table 3. The full SODACER–Sophia framework consistently achieved better performance than RER and CBER across the evaluated metrics. Removing the dual‑buffer structure increases convergence time and final cost, confirming its role in balancing recent and historical data. Disabling adaptive clustering leads to higher memory consumption, highlighting its importance for experience compression. Finally, substituting Sophia with Adam slows convergence, demonstrating the advantage of adaptive second‑order updates.
Table 3: Quantitative ablation study results for the SODACER–Sophia framework
Configuration Final Cost $J$ Convergence Steps Memory (MB)
Full SODACER–Sophia 1.00 15,000 45
Single‑buffer only 1.18 18,200 50
Without clustering 1.12 16,500 75
With Adam optimizer 1.05 18,800 45
@@
Figure 5: Spectral representation of control signals using the proposed approach, based on 200 simulation runs. Optimal control signals exhibit non-zero steady-state values due to the endemic nature of the disease and the cost structure of the intervention problem.
@@
Additionally, to evaluate the effectiveness of control strategies for the HPV model, five distinct cost functions were formulated, each representing a specific control scenario. In “Scenario 1" ($f_{1}$), the vaccination rates of non–sexually active individuals ($w_{1},w_{2}$) were activated, while the vaccination rates of sexually active individuals ($u_{1},u_{2}$) and the female screening rate ($\alpha$) were set to zero. This scenario isolates the effect of early vaccination prior to sexual activity. “Scenario 2" focuses on vaccination of pre-sexually active individuals, utilizing only $w_{1}(t)$ and $w_{2}(t)$ to assess the impact of early immunization on reducing HPV prevalence. “Scenario 3" targets vaccination of sexually active individuals by employing $u_{1}(t)$ and $u_{2}(t)$, highlighting the role of immunizing sexually active populations in controlling transmission. “Scenario 4" isolates the screening rate $\alpha(t)$ as the sole control, emphasizing the effectiveness of detection and early intervention strategies. Lastly, in “Scenario 5", all controls ($w_{1}(t),w_{2}(t),u_{1}(t),u_{2}(t),\alpha(t)$) are set to fixed values, providing a baseline for evaluating the dynamic optimization capabilities of the proposed methodology. These scenarios enable a systematic analysis of individual and combined control strategies to optimize HPV intervention. Table 4 presents a summary of scenarios, including their corresponding objective functions and activated control variables.
Table 5 presents the comparative performance of RER, CBER, and SODACER across the five control scenarios ($f_{1}$$f_{5}$). The numerical values for each method and scenario are the mean final cost achieved over 200 independent runs conducted under identical experimental conditions, including randomized initial states and consistent hyperparameters. The Friedman Rank row displays the average rank (where 1 is best and 3 is worst) for each method across all scenarios, providing a non‑parametric statistical summary of overall performance.
@@
The performance results, summarized in Table 5, demonstrate that SODACER-Sophia consistently outperformed RER and CBER across all metrics. For instance, in Scenario 5, in which the full set of control variables is included in the system and none of them are set to zero, SODACER‑Sophia achieved the most favorable convergence rate and sample efficiency among the evaluated methods. Figure 6 further illustrates the average results across 200 runs, showing SODACER-Sophia’s improved convergence performance compared to RER and CBER. Additionally, Figure 7 provides a detailed comparison of the performance spectrum across 200 runs for the value of the cost function in Equation (28). The results reveal that SODACER-Sophia exhibited the narrowest gap between the worst and best cases, indicating robust and reliable performance.
@@
Statistical Analysis of Performance Robustness
To quantitatively substantiate claims regarding variance reduction and algorithmic robustness, we provide a detailed statistical analysis of the results obtained from the 200 independent runs. Table LABEL:tab:uncertainty_metrics reports, for each method and each control scenario ($f_{1}$$f_{5}$), the following key uncertainty measures: the standard deviation $\sigma$ (a direct measure of dispersion), the variance $\sigma^{2}$, the coefficient of variation $CV=\sigma/\mu$ (a normalized measure of variability), and the 95% confidence interval. The confidence interval is computed as $\mu\pm 1.96\,\sigma/\sqrt{200}$, quantifying the precision of the performance estimate.
The data reveal a consistent and significant pattern: the proposed SODACER–Sophia framework exhibits the smallest standard deviation and variance across all scenarios. For instance, in Scenario $f_{5}$, SODACER–Sophia attains a standard deviation of 0.09, which is an order of magnitude smaller than that of RER (1.05) and three times smaller than that of CBER (0.31). This demonstrates a markedly more stable and predictable learning process.
The coefficient of variation (CV) provides a scale-invariant metric for comparing relative variability across methods with different performance baselines. A lower CV indicates greater robustness relative to the achieved performance level. SODACER–Sophia consistently yields the lowest CV values (ranging from 6.6% to 9.0%), confirming its improveed stability. In contrast, the baseline methods show CV values between 12.9% and 19.2%, reflecting higher relative uncertainty in their outcomes.
The narrow 95% confidence intervals for SODACER–Sophia further underscore the reliability of its performance estimates. The tight bounds, particularly evident in challenging scenarios like $f_{5}$, indicate high confidence in its improved performance. Collectively, these statistical metrics provide rigorous, quantitative evidence supporting our claims of reduced variance and enhanced robustness, affirming that the proposed framework delivers not just better but also more dependable optimal control policies.
Table 6: Statistical performance metrics across 200 independent runs
Scenario Method Statistical Metrics
Mean ($\mu$) Std ($\sigma$) Variance ($\sigma^{2}$) CV (%) 95% CI (Lower, Upper)
$f_{1}$ RER 2.84 0.52 0.27 18.3 (2.77, 2.91)
CBER 3.20 0.48 0.23 15.0 (3.13, 3.27)
SODACER 2.73 0.18 0.03 6.6 (2.71, 2.75)
$f_{2}$ RER 2.43 0.45 0.20 18.5 (2.37, 2.49)
CBER 2.07 0.31 0.10 15.0 (2.03, 2.11)
SODACER 1.69 0.12 0.01 7.1 (1.67, 1.71)
$f_{3}$ RER 2.67 0.41 0.17 15.4 (2.61, 2.73)
CBER 2.32 0.35 0.12 15.1 (2.27, 2.37)
SODACER 1.78 0.15 0.02 8.4 (1.76, 1.80)
$f_{4}$ RER 3.87 0.68 0.46 17.6 (3.78, 3.96)
CBER 3.37 0.52 0.27 15.4 (3.30, 3.44)
SODACER 2.89 0.24 0.06 8.3 (2.86, 2.92)
$f_{5}$ RER 5.47 1.05 1.10 19.2 (5.32, 5.62)
CBER 2.40 0.31 0.10 12.9 (2.36, 2.44)
SODACER 1.00 0.09 0.01 9.0 (0.99, 1.01)
Quantitative Safety Performance Metrics Across 200 Independent Runs
The table 7 reports explicit safety metrics for RER, CBER, and SODACER–Sophia across the five control scenarios ($f_{1}$$f_{5}$). Metrics are computed from 200 independent runs. Constraint Violation Rate (CVR) is the percentage of time steps where the state violated the CBF constraint $h(x(t))<0$. Maximum Constraint Violation (MCV) is the worst-case violation magnitude observed. Safe Convergence Percentage (SCP) indicates the percentage of runs where the final policy remained within the safe set throughout evaluation. ASM (Average Safety Margin) is the mean value of $h(x(t))$ across visited states. Higher SCP and ASM, and lower CVR and MCV, indicate superior safety performance.
Table 7: Quantitative safety performance metrics across 200 independent runs
Scenario Method CVR (%) MCV SCP (%) ASM
$f_{1}$ RER 4.50 0.125 87 0.045
CBER 3.20 0.098 92 0.067
SODACER 0.00 0.000 100 0.152
$f_{2}$ RER 5.80 0.142 82 0.038
CBER 3.90 0.105 90 0.071
SODACER 0.00 0.000 100 0.168
$f_{3}$ RER 3.75 0.118 89 0.052
CBER 2.65 0.087 94 0.083
SODACER 0.00 0.000 100 0.181
$f_{4}$ RER 6.25 0.156 78 0.031
CBER 4.40 0.112 88 0.062
SODACER 0.00 0.000 100 0.145
$f_{5}$ RER 8.10 0.183 72 0.024
CBER 5.30 0.134 85 0.058
SODACER 0.00 0.000 100 0.139
Safety Performance Analysis.
The quantitative safety metrics presented in Table 7 provide explicit validation of the safety guarantees claimed by the SODACER–Sophia framework. The results demonstrate that the integration of Control Barrier Functions (CBFs) within the optimization loop ensures strict adherence to state constraints. SODACER–Sophia achieves a perfect 0% constraint violation rate (CVR) and 100% safe convergence (SCP) across all five scenarios and all 200 independent runs, confirming that safety is never compromised during learning or deployment. In contrast, both baseline methods (RER and CBER) exhibit non-negligible violation rates and fail to guarantee safe convergence in a significant portion of runs. Furthermore, the Average Safety Margin (ASM) for SODACER–Sophia is consistently higher, indicating that the system operates further from the constraint boundary, providing a practical buffer against uncertainties and disturbances. These metrics collectively verify that the proposed framework successfully translates the theoretical safety assurances of CBFs into empirically robust safe behavior.